OpenStudy (anonymous):

Can someone help to walk me through the steps for this! Find an equation of the tangent line to the parabola at the given point. x2 = 9y, (–7, 49/9) ______________ Find the x-intercept of the line. (x,y)=(_________) Thanks!

OpenStudy (anonymous):

the equation of the parabola is \[x ^{2}=9y\] ??

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Do you know the equation of the tangent lines of the parabola? If you have this equation \[(x-h)^2=4p(y-k)\] the equation is: \[(x _{0}-h)(x-h)=4p \left[ \left( \frac{ y+y _{0} }{ 2 } -k \right) \right]\] you have one point , so replace in xº and yº (x)(-7) = 9[(y+49/9)]/2 operating.. -14x=9y+49 (-49-14x)/9=y So the slove is -14/9 and the intercept -49/9

OpenStudy (anonymous):

Ok! I kind of get how you solved the equation, but how did you find the x-intercept?

OpenStudy (anonymous):

iam sorry

OpenStudy (anonymous):

i give you the intercept with y

OpenStudy (ranga):

x^2 = 9y y = 1/9 * x^2 Slope = dy/dx = 2/9 * x slope at (–7, 49/9) put x = -7 above slope - 2/9 * (-7) = -14/9 You know the alope of the tangent line and you know one point (–7, 49/9) Find the equation of the line: y = mx + b y = -14/9x + b x = -7, y = 49/9 49/9 = -14/9 * (-7) + b 49/9 = 98/9 + b b = 49/9 - 98/9 = (49-98)/9 = -49/9 y = -14/9x - 49/9 They are asking for the x-intercept (NOT the y-intercept). To find the x-intercept set y = 0 and solve for x 0 = -14/9x - 49/9 14/9x = 49/9 x = 49/9 * 9/14 = 49/14 x intercept = 49/14

OpenStudy (anonymous):

if you want the intercept of x , put 0 in y

OpenStudy (anonymous):

For some reason, it keeps getting marked wrong! But the equation is right and I'm starting to understand the concept a little bit better. Thank you so much you two!

OpenStudy (ranga):

oops. sign mistake. x-intercept = -49/14

OpenStudy (anonymous):

That worked :) Thank you!

OpenStudy (ranga):

you are welcome.